OpenStudy (anonymous):

if f(x) = x+2 and g(x) = 3x find f(x)/G(x)

OpenStudy (astrophysics):

You're just dividing, \[\frac{ f(x) }{ g(x) } = \frac{ x+2 }{ 3x }\] does that make sense?

OpenStudy (anonymous):

lol yeah, you are not even really dividing just put one on top of the other

OpenStudy (astrophysics):

Haha yeah xD

OpenStudy (anonymous):

so I'm basically just switching the numbers around lol ? how do I know which order to go in though ?

OpenStudy (astrophysics):

Naw you're not even dividing as sat has pointed out, notice you want \[\frac{ f(x) }{ g(x) }\] so you're basically substituting the values \[f(x) = x+2\] and \[g(x) = 3x\] so putting it together \[\frac{ f(x) }{ g(x) } \implies \frac{ x+2 }{ 3x }\]we have

OpenStudy (anonymous):

Ok I get it a little bit more now that I can see is this problem basically the same thing If f(x) = 3x + 5 and g(x) = 2x - 9, find f(x) - g(x). 5x + 4 x - 4 x + 14 5x - 4

OpenStudy (astrophysics):

Yeah, but you'll have to simplify this one once you plug it in

OpenStudy (astrophysics):

So how would this one look?

OpenStudy (anonymous):

the answer would be 5x - 4 because 3x + 2x = 5x and 5 - 9 equals -4 equaling 5x - 4

OpenStudy (astrophysics):

\[f(x)-g(x) = 3x+5-(2x-9)\] we must be careful

OpenStudy (astrophysics):

Note the brackets

OpenStudy (anonymous):

ok so I was wrong ? and I know I always forget to use the brackets .

OpenStudy (astrophysics):

Yeah because we are subtracting by all of g(x) so we must distribute the negative over 2x and -9 so we get \[f(x)-g(x) = 3x+5-(2x-9) = 3x+5-2x+9\]

OpenStudy (astrophysics):

Now simplify and you're done

OpenStudy (anonymous):

so 5x - 4 would be my answer ? or x + 14 ?

OpenStudy (astrophysics):

3x-2x = x 5+9 = 14 So we have x+14

OpenStudy (anonymous):

The inverse of f(x) = x + 2 is f-1(x) = 2x − 4. TRUE FALSE I'm sorry if I'm asking to many questions

OpenStudy (astrophysics):

To find the inverse: Replace f(x) with y Switch x's and y's, so put x where y is and y where x is. Solve for y Replace y with f^-1(x)

OpenStudy (anonymous):

so it would look like x= y + 2 is f - (y) = 2y - 4 ?